OpenStudy (anonymous):

Find a rational number and an irrational number that are between 7.7 and 7.9. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

@freckles @confluxepic @pooja195

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@♪Chibiterasu

OpenStudy (anonymous):

@Michele_Laino can you please post a summarry of the answer??

OpenStudy (anonymous):

we already did this one its just that i dont know what to put sorry

OpenStudy (anonymous):

@Michele_Laino are you done?

OpenStudy (michele_laino):

first part: we have to search for a rational number q, such that: 7.7< q < 7.9 Now I note that: 7.7=77/10, and 7.9=79/10, so I have to search for a rational number q such that the equivalent condition holds: \[\Large \frac{{77}}{{10}} < q < \frac{{79}}{{10}}\] a possible choice for q, is: \[\Large q = \frac{{78}}{{10}} = 7.8\]

OpenStudy (anonymous):

oh sorry ok so thats the answer to the question??

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

it is the answer to the first part, since we have to find an irrational number